# On manifolds with almost non-negative Ricci curvature and integrally-positive $k^{th}$-scalar curvature

### v2

@@ Abstract
If $(M^{n},g)$ is a Riemannian manifold satisfying such bounds for $k\geq 2$ and additionally the Ricci curvature is asymptotically non-negative, then we show that $M$ has at most $(k-1)$-dimensional behavior at large scales.
If $k=n={\rm dim}(M)$, so that the integral lower bound is on the scalar curvature, assuming in addition that the $(n-2)$-Ricci curvature is asymptotically non-negative, then we prove that the dimension drop at large scales improves to $n-2$.
@@ 1 Introduction
where $\mathsf{R}$ denotes the scalar curvature. Under the assumption of non-negative sectional curvature, the aforementioned bound on the scalar curvature was obtained by Petrunin in [PetruninYau]. A non-collapsed version of Yau’s problem was then proposed by Naber in [Naberconjectures], where he conjectured that if $(M^{n},g,p)$ is a pointed Riemannian manifold with $\mathsf{Ric}_{M}\geq-\lambda$ and $\mathsf{Vol}(B_{1}(p))\geq v$, then $\fint_{B_{1}(p)}\mathsf{R}\,{\rm d}\mathsf{Vol}\leq C(\lambda,n,v)$. If true, Naber’s version of Yau conjecture would allow to define a scalar curvature measure on non-collapsed Ricci limit spaces (see again [Naberconjectures]).
These conjectures are still open and, more generally, the geometric interplay between Ricci and scalar curvature bounds is still not fully understood. A step forward in this direction was carried out in a series of recent papers showing that $3$-manifolds with non-negative Ricci curvature and scalar curvature bounded from below by a strictly positive constant have at most linear volume growth at infinity (see [wangvolumethreemanifolds, Chodoshvolumethreemanifolds]). This is a particular case of a more general conjecture by Gromov [LargeGromov] stating that a non-compact pointed Riemannian manifold $(M^{n},g,p)$ of dimension $n$ with non-negative Ricci curvature and scalar curvature bounded below by a strictly positive constant should satisfy
@@
Without the condition on the Ricci curvature, manifolds with uniformly positive scalar curvature can have arbitrarily large volume growth (by the surgery results in [GromovLawsonSurgery1, GromovLawsonSurgery2]). Nevertheless such $3$-manifolds have been shown to have bounded Urysohn $1$-width (see [Katzscalar, GromovScalar, Liokumovitchscalar, liokumovitchwang] and, assuming also non-negative Ricci curvature, [ZhuGeometry]). This class of results enters in the framework of another more general conjecture by Gromov [GromovScalar], stating that a manifold of dimension $n$ with scalar curvature bounded below by a strictly positive constant should have bounded Urysohn $(n-2)$-width.
@@
Given a manifold $(M^{n},g)$ and $k\in\mathbb{N}$, $k\leq n$, we denote by $\mathsf{R}_{k}$ the sum of the lowest $k$ eigenvalues, counted with their multiplicity, of the Ricci tensor $\mathsf{Ric}_{M}$. In analogy with the terminology used in [ShenKricci], if $\mathsf{R}_{k}\geq c$, we say that the $k^{th}$-scalar curvature of $M$ is bounded below by $c$ (note that this is different from the convention used in [WolfsonKricci], where the Ricci curvature is said to be $k$-positive if $\mathsf{R}_{k}\geq 0$). In this paper, we study the geometry of manifolds with almost non-negative Ricci curvature and $k^{th}$-scalar curvature bounded from below in integral sense by a positive constant (see [Moull1, Moull2, Moull3, Moull4, Moull5] for previous literature on manifolds with $\mathsf{R}_{k}\geq c$). As a byproduct, we also obtain new results on manifolds with almost non-negative Ricci curvature and a positive lower integral bound on the scalar curvature.
Broadly speaking, we show that a lower Ricci curvature bound combined with a positive lower bound on the $k^{th}$-scalar curvature forces the manifold to have $(k-1)$-dimensional behavior at large scales. Moreover, if $k$ coincides with the dimension of the manifold (so that the lower bound is on the scalar curvature), the dimensional bound on the geometry at large scales can be improved to $(k-2)$, matching in this way the dimension-drop appearing in the aforementioned conjectures by Gromov. Our approach relies on techniques from metric geometry in which a limit space arises from a sequence of manifolds (see the foundational papers [ChCo1, Colding1, ChcoStructure2, ChCo3]).
The first main result of the paper is Theorem 1.1, which shows the continuity of certain integral functionals of $\mathsf{R}_{k}$ under Gromov-Hausdorff convergence. We recall that, if Yau conjecture holds, given a sequence $(M_{i},g_{i},p_{i})\to(X,\mathsf{d},x)$, one can define a scalar curvature measure $\mu$ on $X$ as a weak limit $\mathsf{R}_{M_{i}}\mathsf{Vol}_{i}\to\mu$ (see [Naberconjectures, Conjecture 2.18]). Hence, continuity results for $\int\mathsf{R}$ under GH convergence would ensure that, if $X$ is itself a smooth manifold, then $\mu=\mathsf{R}_{X}\mathsf{Vol}$. Theorem 1.1 proves a result along these lines when $X=\mathbb{R}^{n}$ (see Remark 3.5). Other continuity results for the integral of the scalar curvature were previously obtained in the presence of a lower bound on the sectional curvature in [PetruninPoli, Theorem B3] (after Perelman) and [lebedeva2022curvaturetensorsmoothablealexandrov]. We refer to Definition 2.5 for the definition of $(k,\delta)$-symmetric and $\delta$-regular balls in a manifold. Given $a,b\in\mathbb{R}$, we denote $a\wedge b:=\min\{a,b\}$ and $a\vee b:=\max\{a,b\}$.
@@
[…]
converge to 00 as $\delta\searrow 0$.
@@
Theorem 1.2 has far reaching implications both for the metric and the topological properties of manifolds satisfying (i), (ii), or (iii). These are collected in Theorem 1.3 below. We denote by $\mathrm{b}_{1}(M)$ the first Betti number of $M$. Given a loop $\gamma$ in $M$, we denote by $[\gamma]$ its equivalence class in the fundamental group $\pi_{1}(M)$. We recall that a metric space $(X,\mathsf{d})$ has $1$-Urysohn width $\leq d$ if there exists a $1$-simplex $Y$, and a continuous map $\pi:X\to Y$, such that $\mathrm{diam}(f^{-1}(y))\leq d$ for every $y\in Y$.
@@
Let $v,\epsilon,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. There exist $C,\delta>0$ with the following property. Let $(M^{n},g,p)$ be a manifold satisfying one of the conditions (i), (ii), (iii). Then,
@@
    If there exists a loop $\gamma\subset M$ such that $[\gamma]\in\pi_{1}(M)$ has infinite order, then $M$ is contained in a $C$-neighbourhood of $\gamma$.
The previous two theorems are part of a series of recent results where positive lower bounds on the (scalar) curvature imply $1$-dimensional behavior at large scales both from the metric and the topological viewpoint. For instance, a linear volume growth bound for $3$-manifolds with $\mathsf{Ric}\geq 0$ and $\mathsf{R}\geq 1$ was obtained in [wangvolumethreemanifolds], and was then generalized in [Chodoshvolumethreemanifolds, antonelli2024newspectralbishopgromovbonnetmyers, zhou2024optimalvolumeboundvolume]. Similarly, $3$-manifolds with $\mathsf{R}\geq 1$ have been recently shown to have bounded Urysohn $1$-width (see [Katzscalar, GromovScalar, Liokumovitchscalar, liokumovitchwang]).
We highlight that an analog of Theorem 1.2 is not to be expected in the higher dimensional case (see Remark 5.9) and that the result fails without the lower bounds on the Ricci curvature (see Remark 5.7). The proof of Theorem 1.2 relies on the study of a geometric condition on length minimizing geodesics in a metric space that was first studied in [TrianglesKapo] (see Section 4).
We now turn our attention to manifolds with an integral lower bound on $\mathsf{R}_{k}$, possibly for $k\geq 3$. Theorem 1.4 shows that asymptotically non-negative Ricci curvature (see Remark 3.9), coupled with a sufficiently slow asymptotic integral decay of $\mathsf{R}_{k}$, implies a dimension-drop for the tangent cone at infinity. An analogous result (under different assumptions) was recently proved in [zhu2024twodimensionvanishingsplittingpositive, Theorem 1.6]. We recall that if a pointed non-compact manifold $(M^{n},g,p)$ has non-negative Ricci curvature, then Gromov’s pre-compactness theorem ensures that, for every sequence $(M^{n},g/r_{j},p)$ as $r_{j}\to+\infty$, there exists a subsequence converging in pointed Gromov-Hausdorff sense to a metric space $(X,\mathsf{d},p_{\infty})$. Such space obtained via blow-down is called a tangent cone at infinity of $M$. A similar procedure is possible when the manifold has only asymptotically non-negative Ricci curvature (see again Remark 3.9). Finally, we refer to Section 3.2 for the definition of the $(n-2)^{th}$-Ricci curvature $\mathsf{Ric}_{n-2}$ (after [ShenKricci]).
@@
Let $(M^{n},g,p)$ be a Riemannian manifold a let $L>0,\;\alpha\in(0,2)$.
@@
    Assume that
    $\liminf_{x\to\infty}\mathsf{d}(p,x)^{2}\mathsf{Ric}(x)\wedge 0=0\quad\text{and}\quad\lim_{r\to+\infty}\fint_{B_{r}(p)}0\vee(r^{2-\alpha}\mathsf{R}_{k})\wedge L\,{\rm d}\mathsf{Vol}=L.$
    Let $(X,\mathsf{d},x_{\infty})$ be any tangent cone at infinity of $M$. If $(\mathbb{R}^{d},\mathsf{d}_{eu})$ is a tangent space to $X$ at a point $x\neq x_{\infty}$, then $d\leq k-1$.
@@
    Assume that
    $\liminf_{x\to\infty}\mathsf{d}(p,x)^{2}\mathsf{Ric}_{n-2}(x)\wedge 0=0\quad\text{and}\quad\lim_{r\to+\infty}\fint_{B_{r}(p)}0\vee(r^{2-\alpha}\mathsf{R})\wedge L\,{\rm d}\mathsf{Vol}=L.$
    Let $(X,\mathsf{d},x_{\infty})$ be any tangent cone at infinity of $M$. If $(\mathbb{R}^{d},\mathsf{d}_{eu})$ is a tangent space to $X$ at a point $x\neq x_{\infty}$, then $d\leq n-2$.
The previous theorem is sharp in the sense that it fails for $\alpha=0$ (see Remark 3.16). A biproduct of Theorem 1.4 is that, if a manifold has asymptotically non-negative Ricci curvature and sufficiently slow integral asymptotic decay of the Ricci tensor, then it is compact (see Remark 3.17).
We now consider the topological implications of almost non-negative Ricci curvature and positive $k^{th}$-scalar curvature. This is the content of Theorems 1.5 and 1.6 (compare with [zhu2024twodimensionvanishingsplittingpositive, Theorems 1.3 and 1.7], after [PanBlowDown]).
@@
The proof of Theorem 1.5 follows the lines of [zhu2024twodimensionvanishingsplittingpositive, Theorem 1.3] by replacing the use of [zhu2024twodimensionvanishingsplittingpositive, Theorem 1.6] with Theorem 1.4. We remark that Theorem 1.5, unlike [zhu2024twodimensionvanishingsplittingpositive, Theorem 1.3], allows for the scalar curvature to converge to zero at infinity.
@@
If the manifold in the previous theorem has non-negative Ricci curvature, the previous result follows immediately from the splitting theorem [SplittingCheegerGromoll]. We remark that Theorem 1.6, unlike [zhu2024twodimensionvanishingsplittingpositive, Theorem 1.7], only requires a local bound on the scalar curvature.
Finally, we mention that Theorem 1.1 can be combined with the results of [CheegerNaber] to prove a result from Jiang and Naber (see Theorem 1.7 below). This result was announced in [Naberconjectures] without proof; we sketch one in the appendix.
@@
$\fint_{B_{1}(x)}|\mathsf{Ric}|^{s}\,{\rm d}\mathsf{Vol}\leq C(K,n,v,s).$
@@ Acknowledgments
A. C. wishes to thank Xingyu Zhu for useful discussions on the content of this paper and, in particular, for pointing out the reference [zhu2024twodimensionvanishingsplittingpositive], whose content led to Theorem 1.5.
@@ 2 Preliminaries
We start by recalling some basic notation and terminology about metric spaces. Throughout the section, $(X,\mathsf{d})$ denotes a proper, complete, separable, length metric space. A metric measure spaces $(X,\mathsf{d},\mathfrak{m})$ is a triple, where $(X,\mathsf{d})$ is a metric space and $\mathfrak{m}$ is a non-negative Borel measure on $X$ which is finite on bounded sets and whose support is the whole $X$. We denote by $\mathsf{H}^{k}$ the $k$-dimensional Hausdorff measure induced by $\mathsf{d}$ on $X$. Given a smooth manifold $(M,g)$ we denote by $\mathsf{d}$ and $\mathsf{Vol}$ respectively the Riemannian distance and volume measure induced by $g$. Given an open set $\Omega\subset X$, we denote by $\mathsf{Lip}(\Omega)$, $\mathsf{Lip}_{\rm loc}(\Omega)$ and $\mathsf{Lip}_{c}(\Omega)$ respectively the set of Lipschitz, locally Lipschitz and compactly supported Lipschitz functions in $\Omega$. We denote the slope of $f\in\mathsf{Lip}_{\rm loc}(\Omega)$ at $x\in\Omega$ by
@@
We next recall the main definitions concerning the pointed Gromov-Hausdorff convergence, referring to [YuBurago_1992], [Vil], and [GMS13] for an overview on the subject.
@@
The previous definition of pointed Gromov Hausdorff convergence is slightly different from the one given in [YuBurago_1992, p. 272]. Nevertheless, they coincide when considering the convergence of length spaces (as it will be, throughout the paper).
@@
Both pointed Gromov-Hausdorff and pointed measured Gromov-Hausdorff convergence are metrizable. The corresponding distances are denoted by $\mathsf{d}_{{\rm pGH}}$ and $\mathsf{d}_{{\rm pmGH}}$. We recall that, in the case of a sequence of uniformly locally doubling metric measure spaces $(X_{j},\mathsf{d}_{j},\mathfrak{m}_{j},\bar{x}_{j})$ (as in the case of $\mathsf{RCD}(K,N)$ spaces), pointed measured Gromov-Hausdorff convergence to $(X,\mathsf{d},\mathfrak{m},x)$ can be equivalently characterized by asking for the existence of a proper metric space $(Z,\mathsf{d}_{z})$ such that all the metric spaces $(X_{j},\mathsf{d}_{j})$ are isometrically embedded into $(Z,\mathsf{d}_{z})$, $\bar{x}_{j}\to x$ and $\mathfrak{m}_{j}\to\mathfrak{m}$ weakly in duality with continuous boundedly supported functions in $Z$ (see [GMS13]).
Next, we recall the fundamental notion of $(k,\delta)$-splitting map (see [ColdingConv, ChCo1, Colding1, CheegerNaberCodim4]) on a Riemannian manifold with Ricci curvature bounded below.
@@
The next theorem corresponds to [CJN, Theorem $4.11$] (after [ChCo1] and [Colding1]).
@@
The next result was proved in [ChCo1] (see also [AMSbakry]).
@@
The next theorem follows from [Colding1, ColdingConv, ChCo1].
@@
For every $\epsilon>0$ there exists $\delta>0$ such that if $(M^{n},g)$ has $\mathsf{Ric}_{M}\geq-\delta$ and $B_{10}(x)\subset M$ is $\delta$-regular, then $B_{r}(y)$ is $\epsilon$-regular for every $r\in(0,5)$ and every $y\in B_{5}(x)$.
@@
[…]
@@
The next theorem was proved in [ColdingConv, Colding1].
@@
[…]
@@
    Or: $\lim_{j\to+\infty}\mathsf{Vol}_{j}(B_{1}(p_{j}))=0.$ In this case the Hausdorff dimension of $X$ is bounded above by $n-1$.
@@
Let $(X,\mathsf{d},\mathfrak{m},\bar{x})$ be a Ricci limit space. A metric space $(Y,\mathsf{d}_{Y},y)$ is a tangent space at $\bar{x}$ if there exists a sequence $01>r_{j}>$ 0, $r_{j}\searrow 0$ that satisfies
@@
Gromov pre-compactness theorem implies that the set of tangent spaces at a point is always non-empty. Moreover, in [Colding1] it was shown that for non-collapsed Ricci limit spaces all tangent spaces are metric cones.
@@
${\mathcal{S}}_{k}=\left\{x\in X:\ \mbox{no tangent space is isometric to }\mathbb{R}^{k+1}\times Y\mbox{, for any }(Y,\mathsf{d}_{Y})\right\}.$
The next theorem was proved in [Colding1].
@@
We next recall some properties of $\mathsf{RCD}$ spaces and of finite perimeter sets thereof, which will be useful later in the paper. We assume the reader to be familiar with the subject and we refer to the surveys [Asurv, SSurv, GSurv] for an account of the theory. Let us just recall that an $\mathsf{RCD}(K,N)$ space is a metric measure space $(X,\mathsf{d},\mathfrak{m})$ with Ricci curvature bounded below by $K\in\mathbb{R}$ and dimension bounded above by $\,N\in[1,\infty]$ in a synthetic-sense, and whose Sobolev space $W^{1,2}$ is Hilbert. For the present work, it will suffice to consider the finite dimensional case; thus, if not otherwise specified, $N\geq 1$ will be a real number. Recall that such a class is stable under pmGH convergence, it contains Ricci limit spaces and finite dimensional Alexandrov spaces.
The next result follows from [MondinoNaber, BScon] (after the seminal works for Ricci limits [Colding1, ChCo3, ColdingNaber]).

### v3

@@ Abstract
If $(M^{n},g)$ is a Riemannian manifold satisfying such bounds for $k\geq 2$, then we show that $M$ has at most (k1)(k-1)-dimensional behavior at large scales.
If $k=n={\rm dim}(M)$, so that the integral lower bound is on the scalar curvature, assuming in addition that the (n2)(n-2)-Ricci curvature is non-negative, we prove that the dimension drop at large scales improves to n2n-2.
@@ 1 Introduction
where $\mathsf{R}$ denotes the scalar curvature. Under the assumption of non-negative sectional curvature, the aforementioned bound on the scalar curvature was obtained by Petrunin in [48]. A non-collapsed version of Yau’s problem was then proposed by Naber in [45], where he conjectured that if $(M^{n},g,p)$ is a pointed Riemannian manifold with $\mathsf{Ric}_{M}\geq-\lambda$ and $\mathsf{Vol}(B_{1}(p))\geq v$, then $\fint_{B_{1}(p)}\mathsf{R}\,{\rm d}\mathsf{Vol}\leq C(\lambda,n,v)$. If true, Naber’s version of Yau conjecture would allow to define a scalar curvature measure on non-collapsed Ricci limit spaces (see again [45]).
These conjectures are still open and, more generally, the geometric interplay between Ricci and scalar curvature bounds is still not fully understood. A step forward in this direction was carried out in a series of recent papers showing that $3$-manifolds with non-negative Ricci curvature and scalar curvature bounded from below by a strictly positive constant have at most linear volume growth at infinity (see [44, 17]). This is a particular case of a more general conjecture by Gromov [31] stating that a non-compact pointed Riemannian manifold $(M^{n},g,p)$ of dimension $n$ with non-negative Ricci curvature and scalar curvature bounded below by a strictly positive constant should satisfy
@@
Without the condition on the Ricci curvature, manifolds with uniformly positive scalar curvature can have arbitrarily large volume growth (by the surgery results in [29, 28]). Nevertheless such $3$-manifolds have been shown to have bounded Urysohn $1$-width (see [36, 32, 40, 41] and, assuming also non-negative Ricci curvature, [62]). This class of results enters in the framework of another more general conjecture by Gromov [32], stating that a manifold of dimension $n$ with scalar curvature bounded below by a strictly positive constant should have bounded Urysohn (n2)(n-2)-width.
@@
Given a manifold $(M^{n},g)$ and $k\in\mathbb{N}$, $k\leq n$, we denote by $\mathsf{R}_{k}$ the sum of the lowest $k$ eigenvalues, counted with their multiplicity, of the Ricci tensor $\mathsf{Ric}_{M}$. In analogy with the terminology used in [50], if $\mathsf{R}_{k}\geq c$, we say that the $k^{th}$-scalar curvature of $M$ is bounded below by $c$ (note that this is different from the convention used in [60], where the Ricci curvature is said to be $k$-positive if $\mathsf{R}_{k}\geq 0$). In this paper, we study the geometry of manifolds with almost non-negative Ricci curvature and $k^{th}$-scalar curvature bounded from below in integral sense by a positive constant (see [20, 38, 56, 22, 26] for previous literature on manifolds with $\mathsf{R}_{k}\geq c$). As a byproduct, we also obtain new results on manifolds with almost non-negative Ricci curvature and a positive lower integral bound on the scalar curvature.
Broadly speaking, we show that a lower Ricci curvature bound combined with a positive lower bound on the $k^{th}$-scalar curvature forces the manifold to have (k1)(k-1)-dimensional behavior at large scales. Moreover, if $k$ coincides with the dimension of the manifold (so that the lower bound is on the scalar curvature), the dimensional bound on the geometry at large scales can be improved to (k2)(k-2), matching in this way the dimension-drop appearing in the aforementioned conjectures by Gromov. Our approach relies on techniques from metric geometry in which a limit space arises from a sequence of manifolds (see the foundational papers [8, 9, 10, 11]).
The first main result of the paper is Theorem 1.1, which shows the continuity of certain integral functionals of $\mathsf{R}_{k}$ under Gromov-Hausdorff convergence. We recall that, if Yau conjecture holds, given a sequence $(M_{i},g_{i},p_{i})\to(X,\mathsf{d},x)$, one can define a scalar curvature measure $\mu$ on $X$ as a weak limit $\mathsf{R}_{M_{i}}\mathsf{Vol}_{i}\to\mu$ (see [45, Conjecture 2.18]). Hence, continuity results for $\int\mathsf{R}$ under GH convergence would ensure that, if $X$ is itself a smooth manifold, then $\mu=\mathsf{R}_{X}\mathsf{Vol}$. Theorem 1.1 proves a result along these lines when $X=\mathbb{R}^{n}$ (see Remark 3.5). Other continuity results for the integral of the scalar curvature were previously obtained in the presence of a lower bound on the sectional curvature in [47, Theorem B3] (after Perelman) and [39]. We refer to Definition 2.5 for the definition of $(k,\delta)$-symmetric and $\delta$-regular balls in a manifold. Given $a,b\in\mathbb{R}$, we denote $a\wedge b:=\min\{a,b\}$ and $a\vee b:=\max\{a,b\}$.
@@
$\sup\Big\{\fint_{B_{1}(x)}\mathsf{R}_{k}\wedge L\,{\rm d}\mathsf{Vol}:(M,g,x)\text{ has }\mathsf{Ric}_{M}\geq-\delta,\text{ and }B_{10}(x)\text{ is $(k,\delta)$-symmetric}\Big\},$
$\sup\Big\{\fint_{B_{1}(x)}|\mathsf{R}|^{s}\,{\rm d}\mathsf{Vol}:(M,g,x)\text{ has }\mathsf{Ric}_{M}\geq-\delta,\text{ and }B_{10}(x)\text{ is }\delta\text{-regular}\Big\}$
converge to 0 as $\delta\searrow 0$.
@@
Theorem 1.2 has far reaching implications both for the metric and the topological properties of manifolds satisfying (i), (ii), or (iii). These are collected in Theorem 1.3 below. We denote by $\mathrm{b}_{1}(M)$ the first Betti number of $M$. Given a loop $\gamma$ in $M$, we denote by $\gamma]$ its equivalence class in the fundamental group $\pi_{1}(M)$. We recall that a metric space $(X,\mathsf{d})$ has $1$-Urysohn width $\leq d$ if there exists a $1$-simplex $Y$, and a continuous map $\pi:X\to Y$, such that $\mathrm{diam}(f^{-1}(y))\leq d$ for every $y\in Y$.
@@
Let $v,\epsilon,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. There exist $C,\delta>0$ with the following property. Let $(M^{n},g)$ be a manifold satisfying one of the conditions (i), (ii), (iii). Then,
@@
    If there exists a loop $\gamma\subset M$ such that $\gamma]\in\pi_{1}(M)$ has infinite order, then $M$ is contained in a $C$-neighbourhood of $\gamma$.
The previous two theorems are part of a series of recent results where positive lower bounds on the (scalar) curvature imply $1$-dimensional behavior at large scales both from the metric and the topological viewpoint. For instance, a linear volume growth bound for $3$-manifolds with $\mathsf{Ric}\geq 0$ and $\mathsf{R}\geq 1$ was obtained in [44], and was then generalized in [17, 4, 61]. Similarly, $3$-manifolds with $\mathsf{R}\geq 1$ have been recently shown to have bounded Urysohn $1$-width (see [36, 32, 40, 41]).
We highlight that an analog of Theorem 1.2 is not to be expected in the higher dimensional case (see Remark 5.9) and that the result fails without the lower bounds on the Ricci curvature (see Remark 5.7). The proof of Theorem 1.2 relies on the study of a geometric condition on length minimizing geodesics in a metric space that was first studied in [34] (see Section 4).
We now turn our attention to manifolds with an integral lower bound on $\mathsf{R}_{k}$, possibly for $k\geq 3$. Theorem 1.4 shows that non-negative Ricci curvature, coupled with a sufficiently slow asymptotic integral decay of $\mathsf{R}_{k}$, implies a dimension-drop for the tangent cones at infinity. An analogous result (under different assumptions) was recently proved in [64, Theorem 1.6]. We recall that if a pointed non-compact manifold $(M^{n},g,p)$ has non-negative Ricci curvature, then Gromov’s pre-compactness theorem ensures that, for every sequence $(M^{n},g/r_{j},p)$ as $r_{j}\to+\infty$, there exists a subsequence converging in pointed Gromov-Hausdorff sense to a metric space $(X,\mathsf{d},p_{\infty})$. Such space obtained via blow-down is called a tangent cone at infinity of $M$. Finally, we refer to Section 3.2 for the definition of the $(n-2)^{th}$-Ricci curvature $\mathsf{Ric}_{n-2}$ (after [50]).
@@
Let $(M^{n},g,p)$ be a Riemannian manifold and let $L>0,\;\alpha\in(0,2)$.
@@
    Assume that $\mathsf{Ric}\geq 0$ and
    $\lim_{r\to+\infty}\fint_{B_{r}(p)}0\vee(r^{2-\alpha}\mathsf{R}_{k})\wedge L\,{\rm d}\mathsf{Vol}=L.$
    Let $(X,\mathsf{d},x_{\infty})$ be any tangent cone at infinity of $M$. If $(\mathbb{R}^{d},\mathsf{d}_{eu})$ is a tangent space to $X$ at a point $x\in X$, then $d\leq k-1$.
@@
    Assume that $\mathsf{Ric}_{n-2}\geq 0$ and
    $\lim_{r\to+\infty}\fint_{B_{r}(p)}0\vee(r^{2-\alpha}\mathsf{R})\wedge L\,{\rm d}\mathsf{Vol}=L.$
    Let $(X,\mathsf{d},x_{\infty})$ be any tangent cone at infinity of $M$. If $(\mathbb{R}^{d},\mathsf{d}_{eu})$ is a tangent space to $X$ at a point $x\in X$, then $d\leq n-2$.
The previous theorem is sharp in the sense that it fails for $\alpha=0$ (see Remark 3.11). A biproduct of Theorem 1.4 is that, if a manifold has non-negative Ricci curvature and sufficiently slow integral asymptotic decay of the Ricci tensor, then it is compact (see Remark 3.12).
We now consider the topological implications of non-negative Ricci curvature and asymptotically positive $k^{th}$-scalar curvature. This is the content of Theorems 1.5 and 1.6 (compare with [64, Theorems 1.3 and 1.7], after [46]).
@@
The proof of Theorem 1.5 follows the lines of [64, Theorem 1.3] by replacing the use of [64, Theorem 1.6] with Theorem 1.4. We remark that Theorem 1.5, unlike [64, Theorem 1.3], allows for the scalar curvature to converge to zero at infinity.
@@
If the manifold in the previous theorem has non-negative Ricci curvature, the previous result follows immediately from the splitting theorem [12]. We remark that Theorem 1.6, unlike [64, Theorem 1.7], only requires a local bound on the scalar curvature.
Finally, we mention that Theorem 1.1 can be combined with the results of [15] to prove a result from Jiang and Naber (see Theorem 1.7 below). This result was announced in [45] without proof; we sketch one in the appendix.
@@
$\fint_{B_{1}(p)}|\mathsf{Ric}|^{s}\,{\rm d}\mathsf{Vol}\leq C(K,n,v,s).$
@@ Acknowledgments
The authors wish to thank the anonymous referee for the careful reading and constructive comments that allowed us to improve the manuscript.
@@
A. C. wishes to thank Xingyu Zhu for useful discussions on the content of this paper and, in particular, for pointing out the reference [64], whose content led to Theorem 1.5.
@@ 2 Preliminaries
We start by recalling some basic notation and terminology about metric spaces. Throughout the section, $(X,\mathsf{d})$ denotes a proper, separable, length metric space. A metric measure spaces $(X,\mathsf{d},\mathfrak{m})$ is a triple, where $(X,\mathsf{d})$ is a metric space and $\mathfrak{m}$ is a non-negative Borel measure on $X$ which is finite on bounded sets and whose support is the whole $X$. We denote by $\mathsf{H}^{k}$ the $k$-dimensional Hausdorff measure induced by $\mathsf{d}$ on $X$. Given a smooth manifold $(M,g)$ we denote by $\mathsf{d}$ and $\mathsf{Vol}$ respectively the Riemannian distance and volume measure induced by $g$. Given an open set $\Omega\subset X$, we denote by $\mathsf{Lip}(\Omega)$, $\mathsf{Lip}_{\rm loc}(\Omega)$ and $\mathsf{Lip}_{c}(\Omega)$ respectively the set of Lipschitz, locally Lipschitz and compactly supported Lipschitz functions in $\Omega$. We denote the slope of $f\in\mathsf{Lip}_{\rm loc}(\Omega)$ at $x\in\Omega$ by
@@
We next recall the main definitions concerning the pointed Gromov-Hausdorff convergence, referring to [7], [55], and [23] for an overview on the subject.
@@
The previous definition of pointed Gromov Hausdorff convergence is slightly different from the one given in [7, p. 272]. Nevertheless, they coincide when considering the convergence of length spaces (as it will be, throughout the paper).
@@
Both pointed Gromov-Hausdorff and pointed measured Gromov-Hausdorff convergence are metrizable. The corresponding distances are denoted by $\mathsf{d}_{{\rm pGH}}$ and $\mathsf{d}_{{\rm pmGH}}$. We recall that, in the case of a sequence of uniformly locally doubling metric measure spaces $(X_{j},\mathsf{d}_{j},\mathfrak{m}_{j},\bar{x}_{j})$ (as in the case of $\mathsf{RCD}(K,N)$ spaces), pointed measured Gromov-Hausdorff convergence to $(X,\mathsf{d},\mathfrak{m},x)$ can be equivalently characterized by asking for the existence of a proper metric space $(Z,\mathsf{d}_{z})$ such that all the metric spaces $(X_{j},\mathsf{d}_{j})$ are isometrically embedded into $(Z,\mathsf{d}_{z})$, $\bar{x}_{j}\to x$ and $\mathfrak{m}_{j}\to\mathfrak{m}$ weakly in duality with continuous boundedly supported functions in $Z$ (see [23]).
Next, we recall the fundamental notion of $(k,\delta)$-splitting map (see [18, 8, 9, 16]) on a Riemannian manifold with Ricci curvature bounded below.
@@
The next theorem corresponds to [14, Theorem $4.11$] (after [8] and [9]).
@@
The next result was proved in [8] (see also [1]).
@@
The next theorem follows from [9, 18, 8].
@@
Let $n\in\mathbb{N}$. For every $\epsilon>0$ there exists $\delta=\delta(n,\epsilon)>0$ such that if $(M^{n},g)$ has $\mathsf{Ric}_{M}\geq-\delta$ and $B_{10}(x)\subset M$ is $\delta$-regular, then $B_{r}(y)$ is $\epsilon$-regular for every $r\in(0,5)$ and every $y\in B_{5}(x)$.
@@
$(M^{n}_{j},\mathsf{d}_{g_{j}},\mathsf{Vol}_{j}(B_{1}(p_{j}))^{-1}\,\mathsf{Vol}_{j},p_{j})\longrightarrow(X,\mathsf{d},\mathfrak{m},\bar{x})\quad\text{in pmGH-sense.}$
@@
The next theorem was proved in [18, 9].
@@
    Either: $\limsup_{j\to+\infty}\mathsf{Vol}_{j}(B_{1}(p_{j}))>0.$ In this case the $\limsup$ is a limit and the pGH convergence can be improved to pmGH convergence to $(X,\mathsf{d},\mathsf{H}^{n},\bar{x})$.
@@
    Or: $\lim_{j\to+\infty}\mathsf{Vol}_{j}(B_{1}(p_{j}))=0.$ In this case the Hausdorff dimension of $X$ is bounded above by n1n-1.
@@
Let $(X,\mathsf{d},\mathfrak{m},\bar{x})$ be a Ricci limit space. A metric space $(Y,\mathsf{d}_{Y},y)$ is a tangent space at $\bar{x}$ if there exists a sequence $1>r_{j}>0$, $r_{j}\searrow 0$ that satisfies
@@
Gromov pre-compactness theorem implies that the set of tangent spaces at a point is always non-empty. Moreover, in [9] it was shown that for non-collapsed Ricci limit spaces all tangent spaces are metric cones.
@@
[…]
The next theorem was proved in [9].
@@
We next recall some properties of $\mathsf{RCD}$ spaces and of finite perimeter sets thereof, which will be useful later in the paper. We assume the reader to be familiar with the subject and we refer to the surveys [2, 54, 25] for an account of the theory. Let us just recall that an $\mathsf{RCD}(K,N)$ space is a metric measure space $(X,\mathsf{d},\mathfrak{m})$ with Ricci curvature bounded below by $K\in\mathbb{R}$ and dimension bounded above by $N\in[1,\infty]$ in a synthetic-sense, and whose Sobolev space $W^{1,2}$ is Hilbert. For the present work, it will suffice to consider the finite dimensional case; thus, if not otherwise specified, $N\geq 1$ will be a real number. Recall that such a class is stable under pmGH convergence, it contains Ricci limit spaces and finite dimensional Alexandrov spaces.
The next result follows from [42, 6] (after the seminal works for Ricci limits [9, 11, 19]).